# Beyond Formula Complexity: Effective Information Criterion Improves Performance and Interpretability for Symbolic Regression

### v1

@@ A.2 Enhance Search-based Methods with EIC
  • Genetic programming maintains a population of candidate formulas, where each individual represents a candidate formula. New individuals are generated by crossover or mutation of high-fitness candidates, while low-fitness candidates are eliminated to increase the overall quality of the population. A common fitness function is defined in equation 5, where Complexity measures the structural size of the formula, NMSE denotes the normalized mean squared error ($\text{MSE}/\text{Var}(y)$), and $\eta<1$ is a regularization constant. This formulation penalizes formulas with large errors or excessive complexity, thereby guiding the search toward accurate and compact formulas. It is worth noting that, when calculating the MSE, we first decompose formulas into additive terms and then apply linear regression to fit the coefficients for these terms. This approach integrates the idea of SINDy and effectively improves the performance of algorithms.
  • MCTS constructs a search tree where each node represents a candidate formula. Child nodes are those formulas that can be obtained by mutating parent formulas, and their reward values are computed according to Equation equation 5. In each search iteration, the algorithm starts from the root node and selects a promising leaf node based on its upper confidence bound (UCB) score, which balances average reward and visitation counts. The chosen leaf is then expanded through mutation, and the resulting reward is backpropagated to update the average reward and visitation counts of its ancestors. This process iteratively guides the search toward structurally and numerically promising formulas.
@@
as the modified fitness and reward functions, where $\alpha>0$ penalizes formulas with higher EIC, thereby steering the search away from structurally unreasonable solutions. In this work, we use $\eta=0.999$ in equation 5 as suggested by yu2025symbolic [11] and sun2022symbolic [26]. For the choice of $\alpha$, we note that the first term of equation 5 has a value range of 0 to 1, while, as shown in Figure 2, EIC ranges from 0 to 10. Since EIC serves as an auxiliary search objective, its weight should generally be an order of magnitude smaller than that of the primary objective. Accordingly, we set $\alpha=0.01$ for MCTS. For the genetic algorithm, we observed that it is more sensitive to the auxiliary objective. Therefore, we chose a smaller value of $\alpha=0.002$ to optimally balance the algorithm’s ability to discover formulas with low EIC while still ensuring high $R^{2}$ accuracy.
A.3 Enhance Generative Methods with EIC
To demonstrate that EIC can enhance the sample efficiency and performance of pretraining-based symbolic regression methods, we focus on three representative approaches: E2ESR[7], SNIP[8], and SR4MDL[11], spanning the progression of this research line from earlier efforts to more recent advances in the last year. E2ESR is trained to predict formula tokens directly from $(X,y)$ data pairs; SNIP extends E2ESR with a CLIP-inspired contrastive loss; and SR4MDL further modifies SNIP by changing the prediction target to the minimum description length of formulas to guide the search. All three approaches rely on the formula generation algorithm proposed by lample2019deep [35], which generates formulas with random and diverse forms to pretrain generative models.
@@ A.3 Enhance Generative Methods with EIC
Appendix B Detailed Experimental Results
B.1 Enhance Search-based Methods with EIC
Corresponding to Figures 5 and 6 in the main text, we provide the raw experimental results for 17 baseline methods and our four methods under four noise levels on white-box data and on black-box data, as shown in Tables 3, 4, 5, 6, and 7, respectively. The tables report formula accuracy, complexity, running duration, and the average EIC of the resulting formulas.
@@ B.1 Enhance Search-based Methods with EIC
Type algorithm $R^{2}$ complexity EIC
@@
B.2 Enhance Generative Methods with EIC
We demonstrate that EIC improves the alignment between pre-training formulas and real-world physical formulas by measuring the similarity between randomly generated formulas and EIC-filtered formulas with three datasets of real physical formulas. These datasets include the Feynman dataset, the Strogatz ODE dataset, and the Wiki Named Equation dataset. The first two are derived from the white-box models in SRBench, while the third was collected by guimera2020bayesian [36] from Wikipedia, containing over a thousand named formulas. We cleaned and deduplicated the formulas, removing those with special operators (e.g., matrix operations) and inherently redundant formulas (e.g., numerous occurrences of simple expressions like $a\times b$, retaining only one). The final set contains 940 physical formulas.
@@ B.2 Enhance Generative Methods with EIC
B.3 Expert Evaluation Results
To demonstrate that EIC evaluates formula interpretability in a manner consistent with human intuition, we compared formulas discovered by EIC-MCTS and PySR. These two methods were chosen because, as shown in Section 4.2, they achieved the lowest average EIC scores (EIC-MCTS) and the lowest average complexity (PySR), respectively, while exhibiting comparable $R^{2}$ performance. We focused on the Pareto fronts obtained from 19 one- or two-dimensional black-box problems in SRBench, as these problems are easier for experts to interpret through visual inspection. From each pair of Pareto fronts, we selected formula pairs by minimizing the distance function:
$L(f_{\text{1}},f_{\text{2}})=(\text{Complexity}[f_{\text{1}}]-\text{Complexity}[f_{\text{2}}])^{2}+(R^{2}[f_{\text{1}}]-R^{2}[f_{\text{2}}])^{2}-(\text{EIC}[f_{\text{1}}]-\text{EIC}[f_{\text{2}}])^{2},$ (8)
subject to the following constraints:
@@ B.3 Expert Evaluation Results
where $f_{1}$ and $f_{2}$ traverse each formula on the Pareto front of PySR and EIC-MCTS, respectively. This procedure ensured that selected pairs exhibited similar complexity and accuracy but substantially different EIC values. In total, 172 formula pairs were obtained across the 19 Pareto fronts.
Each pair was then presented to large language models (LLMs) acting as a domain expert to select the preferred one. To provide independent evaluations of interpretability, we employed two large language models: GPT-4o-Mini and Qwen3, which are trained on different linguistic distributions and can thus reduce the risk of model-specific bias. We format the prompt based on the format shown in Figure 13, where we provided the LLms with the dataset name, textual description, $(x,y)$ sampled points, the mathematical expressions of the two formulas, as well as their $R^{2}$ values and complexities. Importantly, the EIC scores were withheld from the models to ensure unbiased evaluation. We evaluate every formula pair on every LLM five times under a temperature setting of 0.7 to account for sampling variability. The result is shown in Figure 14, where LLM shows a preference of 72.19% to formulas discovered by EIC, which is similar to that of the human experts, demonstrating the reliability of the results.
For the human rating experiment, we invite 108 volunteer participants with at least a bachelor’s degree in science or engineering. Each participant was randomly assigned 10 formula pairs. For each pair, participants received the data samples visualization, dataset description, and the two candidate formulas with their $R^{2}$ and complexity, and were asked to choose the more interpretable one (EIC scores were withheld), as demonstrated in Figure 11 and 12. We collected 1080 evaluations in total. After discarding responses completed in under 60 seconds, 840 evaluations remained, with each pair assessed on average 4.9 times.

### v2

@@ B.1 Enhance Search-based Methods with EIC
  • Genetic programming maintains a population of candidate formulas, where each individual represents a candidate formula. New individuals are generated by crossover or mutation of high-fitness candidates, while low-fitness candidates are eliminated to increase the overall quality of the population. A common fitness function is defined as
    $f(C,\text{NMSE};\eta)={\eta^{C}}/{(1+\text{NMSE})},$ (34)
    where Complexity measures the structural size of the formula, NMSE denotes the normalized mean squared error ($\text{MSE}/\text{Var}(y)$), and $\eta<1$ is a regularization constant. This formulation penalizes formulas with large errors or excessive complexity, thereby guiding the search toward accurate and compact formulas. It is worth noting that, when calculating the MSE, we first decompose formulas into additive terms and then apply linear regression to fit the coefficients for these terms. This approach integrates the idea of SINDy and effectively improves the performance of algorithms.
  • MCTS constructs a search tree where each node represents a candidate formula. Child nodes are those formulas that can be obtained by mutating parent formulas, and their reward values are computed according to Equation (34). In each search iteration, the algorithm starts from the root node and selects a promising leaf node based on its upper confidence bound (UCB) score, which balances average reward and visitation counts. The chosen leaf is then expanded through mutation, and the resulting reward is backpropagated to update the average reward and visitation counts of its ancestors. This process iteratively guides the search toward structurally and numerically promising formulas.
@@
as the modified fitness and reward functions, where $\alpha>0$ penalizes formulas with higher EIC, thereby steering the search away from structurally unreasonable solutions. In this work, we use $\eta=0.999$ in (34) as suggested by Yu et al. [34] and Sun et al. [26]. For the choice of $\alpha$, we note that the first term of (34) has a value range of 0 to 1, while, as shown in Figure 2, EIC ranges from 0 to 10. Since EIC serves as an auxiliary search objective, its weight should generally be an order of magnitude smaller than that of the primary objective. Accordingly, we set $\alpha=0.01$ for MCTS. For the genetic algorithm, we observed that it is more sensitive to the auxiliary objective. Therefore, we chose a smaller value of $\alpha=0.002$ to optimally balance the algorithm’s ability to discover formulas with low EIC while still ensuring high $R^{2}$ accuracy.
B.2 Enhance Generative Methods with EIC
To demonstrate that EIC can enhance the sample efficiency and performance of pretraining-based symbolic regression methods, we focus on three representative approaches: E2ESR[12], SNIP[21], and SR4MDL[34], spanning the progression of this research line from earlier efforts to more recent advances in the last year. E2ESR is trained to predict formula tokens directly from $(X,y)$ data pairs; SNIP extends E2ESR with a CLIP-inspired contrastive loss; and SR4MDL further modifies SNIP by changing the prediction target to the minimum description length of formulas to guide the search. All three approaches rely on the formula generation algorithm proposed by Lample and Charton [16], which generates formulas with random and diverse forms to pretrain generative models.
@@ B.2 Enhance Generative Methods with EIC
Appendix C Detailed Experimental Results
C.1 Structural Instability of LLM-Generated Formulas
In addition to search-based and generative symbolic regression methods, we further investigate the structural properties of formulas generated by large language models (LLMs) in LLM-SRBench [25]. This benchmark includes two categories of LLM-generated formulas: Transform and Synth. The Transform subset is obtained by applying symbolic transformations (e.g., inversion) to real physical formulas, while the Synth subset is constructed via LLM-assisted synthesis of known and novel terms, where LLMs combine established scientific components with newly synthesized symbolic terms.
Figure 7 and Figure 8 present the EIC distributions of formulas from SRBench physical formulas, Transform, and Synth subsets, respectively. Both LLM-generated subsets exhibit substantially higher EIC values than real physical formulas from SRBench, indicating a higher degree of structural instability. In particular, formulas in the Synth subset show the highest EIC distribution, suggesting that LLM-synthesized symbolic expressions tend to involve structurally fragile or numerically unstable constructions, despite their surface-level plausibility.
Figure 7: Average EIC comparison between real physical formulas and LLM-generated formulas. Error bars denote 95% confidence intervals.
Figure 8: Violin plots of EIC distributions for SRBench physical formulas, Transform, and Synth subsets. Scattered points indicate outliers.
These results lead to three key observations. First, formulas synthesized by LLMs (Synth) exhibit significantly higher EIC values than real physical formulas, indicating that both LLM-generated symbolic expressions and LLM-based symbolic regression methods inherit the challenge of structural instability. Second, formulas derived from physical transformations (Transform) show notably lower EIC than Synth formulas, providing further evidence that symbolic expressions rooted in natural scientific laws differ fundamentally from those synthesized purely by LLMs. Third, LSR-Transform formulas still exhibit higher EIC than the original SRBench physical formulas. We attribute this increase to the fact that some physical equations are not naturally invertible from a causal or physical perspective. When such equations are algebraically inverted to construct LSR-Transform tasks, the resulting expressions may involve divisions by quantities that can become arbitrarily small over the data domain, thereby introducing numerically and structurally unstable forms.
C.2 Enhance Search-based Methods with EIC
Figure 9 illustrates the distribution of Effective Information Criterion (EIC) values for formulas discovered by all benchmarked methods. The left panel depicts the distributions on white-box problems across varying noise levels. Notably, the integration of EIC consistently drives both MCTS and GP toward lower EIC values, aligning them more closely with the distribution of ground-truth physical formulas compared to their vanilla versions. This confirms that EIC effectively suppresses structurally unreasonable and numerically unstable patterns. The right panel presents the results on black-box problems. Here, incorporating EIC leads to a substantial reduction in the prevalence of high-EIC structures for both MCTS and GP. Crucially, this reduction is more pronounced compared to the white-box setting, suggesting that in the absence of a known underlying physical law, standard SR methods are prone to overfitting via overly complex and unstable functional forms. EIC effectively mitigates this tendency, enforcing structural plausibility even when the ground truth is unknown.
Figure 9: EIC distributions of formulas discovered by baseline methods and our EIC-enhanced variants. The left panel presents results on 133 white-box problems under varying noise levels ($\{0,10^{-3},10^{-2},10^{-1}\}$), while the right panel presents results on 122 black-box problems. The markers represent the mean EIC values across all problems, and error bars indicate 95% confidence intervals. Notably, the EIC-enhanced methods (highlighted in bold color) consistently exhibit lower EIC values compared to their vanilla counterparts, demonstrating that incorporating EIC effectively steers the search toward structurally rational formulas.
Corresponding to Figures 4 (in the main text) as well as Figure 9, we provide the raw experimental results for 17 baseline methods and our four methods under four noise levels on white-box data and on black-box data, as shown in Tables 3, 4, 5, 6, and 7, respectively. The tables report formula accuracy, complexity, running duration, and the average EIC of the resulting formulas.
@@ C.2 Enhance Search-based Methods with EIC
Type algorithm $R^{2}$ $\uparrow$ complexity $\downarrow$ EIC $\downarrow$
@@
C.3 Enhance Generative Methods with EIC
We demonstrate that EIC improves the alignment between pre-training formulas and real-world physical formulas by measuring the similarity between randomly generated formulas and EIC-filtered formulas with three datasets of real physical formulas. These datasets include the Feynman dataset, the Strogatz ODE dataset, and the Wiki Named Equation dataset. The first two are derived from the white-box models in SRBench, while the third was collected by Guimerà et al. [11] from Wikipedia, containing over a thousand named formulas. We cleaned and deduplicated the formulas, removing those with special operators (e.g., matrix operations) and inherently redundant formulas (e.g., numerous occurrences of simple expressions like $a\times b$, retaining only one). The final set contains 940 physical formulas.
@@ C.3 Enhance Generative Methods with EIC
C.4 Expert Evaluation Results
Generation of Formula Pairs for Human Evaluation. To rigorously validate the alignment between EIC and human interpretability preferences, we constructed a dataset of paired formulas by strictly controlling for accuracy and complexity. Specifically, we aggregated formulas discovered by search-based symbolic regression methods in Section 4.1, keeping those one- and two-dimensional black-box problems in SRBench (in total of 19 problems). This resulted in a comprehensive candidate pool covering diverse structural patterns. Then, from this aggregated pool, we selected formula pairs to maximize structural divergence (EIC difference) while maintaining strict performance parity. This was formulated as minimizing the following distance function:
$L(f_{\text{1}},f_{\text{2}})=(\text{Complexity}[f_{\text{1}}]-\text{Complexity}[f_{\text{2}}])^{2}+$ (35)
$(R^{2}[f_{\text{1}}]-R^{2}[f_{\text{2}}])^{2}-(\text{EIC}[f_{\text{1}}]-\text{EIC}[f_{\text{2}}])^{2},$
subject to the following constraints to ensure the condition that all else being equal:
$|\text{Complexity}[f_{\text{1}}]-\text{Complexity}[f_{\text{2}}]|\leq 2,$ (36)
@@ C.4 Expert Evaluation Results
Here, $f_{1}$ and $f_{2}$ represent any two distinct formulas selected from the aggregated pool. The negative sign for the EIC term in $L$ drives the selection toward pairs with the largest possible stability gap, while the constraints ensure that the paired formulas are indistinguishable in terms of fitting accuracy and length. The threshold $R^{2}>0.85$ ensures that only high-quality candidates are evaluated. This procedure yielded a total of 172 valid formula pairs across the 19 problems.
Human Expert Evaluation. For the human rating experiment, we invited 108 volunteer participants with at least a bachelor’s degree in science or engineering. Each participant was randomly assigned 10 formula pairs. For each pair, participants were presented with the data sample visualization, dataset description, and the two candidate formulas alongside their $R^{2}$ and complexity metrics. They were then asked to choose the more interpretable one based on their domain knowledge and intuition (importantly, EIC scores were withheld to ensure blind evaluation), as demonstrated in Figure 13 and 14. We collected 1080 evaluations in total. After discarding responses completed in under 60 seconds to ensure quality, 840 valid evaluations remained, with each pair assessed on average 4.9 times.
LLM Corroboration. To provide independent corroboration of the human evaluation results, we also presented each pair to Large Language Models (LLMs) acting as domain experts. We employed two models, GPT-4o-Mini and Qwen3, which are trained on different linguistic distributions, thereby reducing the risk of model-specific bias. We formatted the prompt as shown in Figure 15, providing the LLMs with the dataset context, sampled points $(x,y)$, mathematical expressions, and performance metrics ($R^{2}$, Complexity), while withholding EIC scores. We evaluated every formula pair on each LLM five times under a temperature setting of 0.7 to account for sampling variability. The results, shown in Figure 12, indicate that LLMs exhibited a preference rate of 72.19% for the lower-EIC formulas. This alignment with human experts (who preferred lower-EIC formulas in 69.8% of cases) further validates the reliability of EIC as a proxy for interpretability.